OpenStudy (anonymous):

(sinx-cosx+1)/(sinx+cosx-1)=sinx+1/cosx

OpenStudy (shubhamsrg):

its a little confusing, can you use more brackets ?

OpenStudy (anonymous):

plz help me to solve me this question (sinx-cosx+1)/sinx+cosx-1)=sinx+1/cosx

OpenStudy (anonymous):

@shivamsharan are you aware of the double angles?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

divide numerator and denominator by cosx..!! and try to proceed..

OpenStudy (anonymous):

bt how plz u show me once i didn't understood please pleease show me once

OpenStudy (anonymous):

sin(2x)=2(sinx)(cosx), and cos(2x)=1-2sin^(2)x=2cos^(2)x-1.

OpenStudy (anonymous):

i didn't get it plz

OpenStudy (anonymous):

well the expression wud be reduced to(tanx-1+secx)/tanx+1-secx).... then use sec^2x-tan^2x=1..then the things will automatically simplify..:)

OpenStudy (anonymous):

how can u show me once by doin it please

OpenStudy (anonymous):

u can't solve it

OpenStudy (anonymous):

no actually it wud become a little bit long,,actually very long...

OpenStudy (anonymous):

k